# Argument Rarity-based Originality Assessment for AI-Assisted Writing

### v1

@@ 5.2 Limitations
Future work should address these limitations and consider the following directions: i) verification of generalizability using diverse datasets, ii) empirical validation through correlation analysis with human evaluation, iii) application to formative assessment through development of feedback generation functionality, and iv) extension to longitudinal studies tracking the development of originality over time. Through these extensions, the proposed framework is expected to contribute further to the evaluation and development of critical thinking skills in educational settings.
@@ 6 Conclusion
As LLMs have become capable of effortlessly generating high-quality text, traditional quality-focused writing assessment is losing its significance. This study proposed AROA, a framework for automatically evaluating argumentative originality in essays. AROA operationally defines originality as rarity within a reference corpus and evaluates it through four complementary components: structural rarity, claim rarity, evidence rarity, and cognitive depth. The framework quantifies the rarity of each component using density estimation and adjusts them with quality scores, thereby treating quality and originality as independent evaluation axes.
Experiments using 1,375 human essays and 1,000 AI essays verified the effectiveness of the proposed framework. A strong negative correlation (r=0.67r=-0.67) was observed between quality and claim rarity, revealing a quality-originality trade-off where higher-quality texts tend to rely on typical claim patterns. Furthermore, while AI essays achieved comparable levels of structural complexity to human essays, their claim rarity was approximately one-fifth that of humans, indicating that LLMs can reproduce the form of argumentation but have limitations in the originality of content.
Future research directions include verification of generalizability across diverse datasets, empirical validation through correlation analysis with human evaluation, and application to formative assessment through development of feedback generation functionality. This study provides a foundation for a paradigm shift in writing assessment in the AI era, from quality to originality.

### v3

@@ 5.4 Limitations and Future Work
Future work should address these limitations through: i) verification of generalizability using diverse datasets, ii) empirical validation through correlation with human expert evaluation, iii) development of feedback generation functionality for formative assessment, and iv) longitudinal studies tracking originality development over time.
@@ 6 Conclusion
This study proposed AROA, a framework for automatically evaluating argumentative originality in essays through four complementary components, structural rarity, claim rarity, evidence rarity, and cognitive depth, using density-based rarity quantification with quality adjustment. Experiments using 1,375 human essays and 1,000 AI essays revealed three key findings: i) a strong quality-originality trade-off (r=0.67r=-0.67 between quality and claim rarity), ii) that contemporary LLMs reproduce structural complexity but exhibit claim rarity at approximately one-fifth of human levels, and iii) that the four components capture genuinely independent aspects of originality. These results demonstrate that quality and originality are independent evaluation axes and that assessment must shift accordingly in the AI era. Future work includes generalizability verification across diverse datasets, empirical validation with human expert evaluation, and development of feedback generation for formative assessment.